In a study of distances traveled by buses before the first major engine failure, a sample of 191 buses resulted in a mean of 96,700 miles and a standard deviation of 37,500 miles. At the 0.05 level of signıficance, test the manufacturer's claim that the mean distance traveled before a major engine failure is more than 90,000 miles.


1. Claim:

Ho:

Ha:


2. Level of Significance:

Test- statistic:

Tails in Distribution:


3. Reject Ho if:


4. Compute for the value of the test statistics.


5. Make a decision:


6. State the conclusion in terms of the original problem.


1
Expert's answer
2021-03-30T07:39:57-0400

1. Claim:

Ho: µ1 = µ0, (the mean distance traveled before a major engine failure is not different from 90,000 miles (µ0))


Ha: µ1 > µ0, (µ0 =90,000 miles), (the mean distance traveled before a major engine failure is more than 90,000 miles


2. Level of Significance: α=0.05


Test- statistic: Z- statistics, "Z=\frac{\bar{X}-\mu_0}{\frac{s}{\sqrt{n}}}"

"Z=\frac{\bar{X}-\mu_0}{\frac{s}{\sqrt{n}}}"

Tails in Distribution: Right tail or upper-tailed test


3. Reject Ho if: Reject H0 if "Z\geq 1.645".


4. Compute for the value of the test statistics.

"Z=" "\frac{\bar{X}-\mu_0}{\frac{s}{\sqrt{n}}}=\frac{96700-90000}{\frac{37500}{\sqrt{191}}}=2.469"

"\frac{\bar{X}-\mu_0}{\frac{s}{\sqrt{n}}}=\frac{96700-90000}{\frac{37500}{\sqrt{191}}}=2.469""\frac{\bar{X}-\mu_0}{\frac{s}{\sqrt{n}}}=\frac{96700-90000}{\frac{37500}{\sqrt{191}}}=2.469"


5. Make a decision: We reject H0 because 2.469 > 1.645.


6. State the conclusion in terms of the original problem.


We have statistically significant evidence at α=0.05, to show that the mean distance traveled before a major engine failure is more than 90,000 miles.